# Wavefunction textures in twisted bilayer graphene from first principles

### v1

@@ Abstract
Discussion.—We have presented a detailed first-principles characterization of the low-energy wavefunctions of magic-angle tBLG on atomic and moiré scales. Both the microscopic and moiré-scale wavefunctions provide valuable insight into important aspects of the low-energy physics of tBLG, including the dependence of the topology on interlayer interactions. Starting from well separated, noninteracting graphene layers, we have demonstrated two distinct band inversions that occur as the interlayer separation decreases, effectively increasing the interlayer interaction. The first occurs when the flat bands exchange character with the neighboring dispersive bands near $\Gamma$: $\Gamma_{0-}\leftrightarrow\Gamma_{1-}$ and $\Gamma_{0+}\leftrightarrow\Gamma_{1+}$. As the interlayer separation is decreased further, the flat bands invert among themselves: $\Gamma_{0-}\leftrightarrow\Gamma_{0+}$. Therefore, this second inversion further entangles bands originating above and below the Dirac point: $\Gamma_{0-}\leftrightarrow\Gamma_{1+}$ and $\Gamma_{0+}\leftrightarrow\Gamma_{1-}$. Thus, minimal phenomenological models must include complementary bands above and below the flat bands to accurately capture the symmetry of the electronic bands of tBLG.
@@
Acknowledgments.—The authors acknowledge support from the Simons Foundation award No. 896626 and the US Army Research Office (ARO) MURI project under Grant No. W911NF-21-0147. A.Z. is supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0025528. M.M.E.A. acknowledges support from the Singapore Ministry of Education, Research Centre of Excellence Institute for Functional Intelligence Materials. E. M. acknowledges the hospitality of the Harvard Physics Department, where part of this work was done. Computing resources were provided by the Harvard University FAS Division of Science Research Computing Group and the Texas Advanced Computing Center (TACC) at The University of Texas at Austin.
Supplemental Material
S1 First-principles calculations
First-principles density functional theory (DFT) calculations were performed using the siesta [53] code, version 5.0, with norm-conserving [63] psml pseudopotentials [64], obtained from Pseudo-Dojo [65]. siesta uses a basis of numerical atomic orbitals (NAOs) [55], and a basis optimized for graphene from Ref. [51] was used in all calculations. In order to keep the calculations efficient, a basis of single-$\zeta$ polarized (SZP) orbitals were used for all calculations, with $2s$, $2p$ orbitals for each carbon atom, and $3d$ polarization orbitals. The optimized basis set made it possible to perform efficient and accurate full first-principles calculations for tBLG for twist angles as low as $\theta=0.99^{\circ}$ (13,468 atoms).
The PBE exchange-correlation functional [66] was used in all calculations, and a DFT-D3 dispersion correction [67] was included in order to treat the long range interactions between the layers. A single $\mathbf{k}$-point was used in all calculations, and the real space grid was determined using an energy cutoff of 300 Ry. Calculations were performed using the MRRR solver for geometry relaxations, including only the first 20 conduction bands above the Fermi level. The SCF was converged until the relative changes in the density matrix were below $10^{-4}$.
Commensurate twisted bilayer structures were generated for twist angles from $\theta=21.79^{\circ}$ down to $\theta=0.99^{\circ}$. Starting with two flat layers with a layer separation of $d=3.35$, the atomic positions were relaxed until the forces on all atoms were below the force tolerance. A force tolerance of 1 meV/Å was used for larger twist angles (smaller supercells), but was increased up to 10 meV/Å for the smaller twist angles (larger supercells) in the interest of computational efficiency. The center of mass was fixed in the out-of-plane direction in order to prevent the bilayer from drifting from its initial position in middle of the cell. After the relaxed geometries were obtained, calculations were performed to obtain the electronic band structures. The default solver was used for these calculations, and all bands were included. The bands were obtained along the path K–$\Gamma$–M–K, with 10 points along each segment.
S2 Wavefunctions
We obtain three-dimensional real-space wavefunctions $\Psi_{n\mathbf{k}}(\mathbf{r})$ for the low-energy bands $n$ at the high-symmetry $\mathbf{k}$ points $\Gamma$, M, and K, in the mBZ. In siesta, atomic-scale wavefunctions are represented as linear combinations of the local (non-orthogonal) basis orbitals $\phi_{\mu}$:
$\Psi_{n\mathbf{k}}(\mathbf{r})=\sum_{\mu}c_{n\mathbf{k},\mu}\phi_{\mu}(\mathbf{r})\;.$ (S1)
Additionally, we obtain sublattice-projected wavefunctions by calculating the projection of $\Psi_{n\mathbf{k}}$ onto the basis orbtials of a specific layer and sublattice $s\in\{A_{1},A_{2},B_{1},B_{2}\}$:
$\Psi_{n\mathbf{k}}^{s}(\mathbf{r})=\sum_{\mu\in s}c_{n\mathbf{k},\mu}\phi_{\mu}(\mathbf{r})\;.$ (S2)
In order to obtain a two-dimensional representation of $\Psi_{n\mathbf{k}}(\mathbf{r})$, we show the modulus squared of the wavefunction, averaged in the out-of-plane direction $z$: ${|\Psi_{n\mathbf{k}}(x,y)|^{2}=\sum_{z}|\Psi_{n\mathbf{k}}(x,y,z)|^{2}}$. For degenerate states, we add the moduli squared of the degenerate wavefunctions. To visualize the moiré-scale features of the atomic-scale wavefunctions, we perform a convolution
$(f*g)(\mathbf{r})=\int f(\mathbf{r}^{\prime})g(\mathbf{r}-\mathbf{r}^{\prime})~{}\mathop{}\!\mathrm{d}\mathbf{r}^{\prime}$ (S3)
of $|\Psi_{n\mathbf{k}}(x,y)|^{2}$ with a symmetric two-dimensional Gaussian filter. The moiré-scale wavefunction $|\widetilde{\Psi}_{n\mathbf{k}}|^{2}$ is given by
$|\widetilde{\Psi}_{n\mathbf{k}}|^{2}(x,y)\equiv(|\Psi_{n\mathbf{k}}|^{2}*\mathcal{N})(x,y)=\sum_{x^{\prime},y^{\prime}}|\Psi_{n\mathbf{k}}(x^{\prime},y^{\prime})|^{2}\mathcal{N}_{x,y}(x^{\prime},y^{\prime})\;.$ (S4)
Here, $\mathcal{N}_{x,y}(x^{\prime},y^{\prime})$ denotes the probability density function of a Gaussian with mean $(x,y)$ and covariance matrix $\sigma^{2}I_{2\times 2}$ evaluated at the point $(x^{\prime},y^{\prime})$. In practice, we sum over all $(x^{\prime},y^{\prime})$ such that $\|(x^{\prime},y^{\prime})-(x,y)\|\leq 3\sigma$. We set $\sigma$ on the order of $1$ to $10$, just large enough to clearly extract the moiré-scale features of the wavefunctions, and generally choose larger $\sigma$ for larger supercells (smaller $\theta$). For example, $\sigma=2.5~{}\mathrm{\AA}$ for $\theta=5.09^{\circ}$ in Fig. 1 in the main text, while $\sigma=10~{}\mathrm{\AA}$ for $\theta=1.08^{\circ}$ in Fig. 2 and 3 in the main text.
Fig. S1 shows the different wavefunction characters mentioned in the main text (AA, AAz, AB/BA, and DW) for $\theta=0.99^{\circ}$. Fig. S2 shows the evolution of the microscopic and moiré-scale wavefunctions at $\Gamma_{0-}$ as a function of twist angle.
Figure S1: Representative wavefunctions for a twist angle of $\theta=0.99^{\circ}$. AA, AAz, AB/BA, and DW wavefunctions are outlined in red, blue, pink, and green, respectively.
Figure S2: Electronic band structures and wavefunctions at $\Gamma_{0-}$ as a function of twist angle $\theta$, with the top row of wavefunction panels showing the microscopic textures and the bottom row showing the moiré-scale features. The color of the circles on the band structures and the borders of the wavefunctions indicates the character of the wavefunctions as described in the main text.
S3 Estimation of Pressure
The energy of magic-angle tBLG is shown as a function of average layer separation in Fig. S3 (a), for a small variation $\Delta z$ about the equilibrium value $z_{f}=3.48$. For each value of $\Delta z$, the atomic coordinates were held fixed and a single point calculation was performed. We note that the minimum in energy is not exactly at $z_{f}$. Due to the large number of degrees of freedom, a very low force tolerance is required to find the true minimum, and the initial relaxation was performed with a force tolerance of 10 meV/Å in the interest of computational efficiency.
An equation of state was used to fit the data and calculate the bulk modulus:
$B=V_{f}\frac{d^{2}E}{dV^{2}}=\frac{z_{f}}{A}\left.\frac{d^{2}E}{dz^{2}}\right|_{z=z_{f}}\;,$ (S5)
where E is the energy, $V=Az$ is the volume, and $A$ is the in-plane area and $z=z_{f}+\Delta z$. A quadratic fit,
$E(z)=\frac{1}{2}V_{f}B(z-z_{0})^{2}\;,$ (S6)
where $z_{0}$ is the energy minimum, is shown in Fig. S3 in red, which yields a bulk modulus of $B=24.95$ GPa. A Birch-Murnaghan equation of state,
$E(z)=\frac{9}{16}V_{f}B\left\{\left[\left(\frac{z_{f}}{z}\right)^{2/3}-1\right]^{3}B^{\prime}+\left[\left(\frac{z_{f}}{z}\right)^{2/3}-1\right]^{2}\left[6-4\left(\frac{z_{f}}{z}\right)^{2/3}\right]\right\}\;,$ (S7)
where $B^{\prime}$ is the derivative of the bulk modulus, is shown in Fig. S3 in blue, yielding a bulk modulus of $B=21.41$ GPa. The Birch-Murnaghan equation of state yields a better fit to the data, as it captures the asymmetry about $\Delta z=0$.
The induced pressure $\Delta P$ as a function of $\Delta z$,
$\Delta P=-\frac{E}{V_{f}}\approx-B\frac{\Delta z}{z_{f}}\;,$ (S8)
is shown in Fig. S3 (b). The solid lines show the linear approximation obtained using the bulk modulus, and the dashed lines show the induced pressure obtained from the derivative of the energy. For the quadratic fit, the two methods match exactly. For the Birch-Murnaghan fit, some nonlinearity is induced. The vertical line shows the approximate $\Delta z/z_{f}$ required to induce the topological phase transition described in the main text. We can see that all estimations indicate a pressure in the range $0.5~{}\mathrm{GPa}<\Delta P<1.0~{}\mathrm{GPa}$ is required to induce this transition.
Figure S3: (a) Energy as a function of average layer separation for magic-angle tBLG. The black circles indicate results from DFT calculations. The red and blue lines show fits using quadratic and Birch-Murnaghan equations of state, respectively. (b) Pressure as a function of layer separation for magic-angle tBLG. The solid lines show the linear approximation using the bulk modulus calculated from the fitting in (a). The dashed lines show the pressure obtained by differentiating the energy. The vertical black line in each plot denotes the value of $\Delta z/z_{f}$ at which the flat bands touch each other at $\Gamma$.

### v2

@@ III Discussion
We have presented a detailed first-principles characterization of the low-energy wavefunctions of magic-angle tBLG on atomic and moiré scales. Both the microscopic and moiré-scale wavefunctions provide valuable insight into important aspects of the low-energy physics of tBLG, including the dependence of the topology on interlayer interactions. Starting from well separated, noninteracting graphene layers, we have demonstrated two distinct band inversions that occur as the interlayer separation decreases, effectively increasing the interlayer interaction. The first occurs when the flat bands exchange character with the neighboring dispersive bands near $\Gamma$: $\Gamma_{0-}\leftrightarrow\Gamma_{1-}$ and $\Gamma_{0+}\leftrightarrow\Gamma_{1+}$. As the interlayer separation is decreased further, the flat bands invert among themselves: $\Gamma_{0-}\leftrightarrow\Gamma_{0+}$. Therefore, this second inversion further entangles bands originating above and below the Dirac point: $\Gamma_{0-}\leftrightarrow\Gamma_{1+}$ and $\Gamma_{0+}\leftrightarrow\Gamma_{1-}$. Thus, minimal phenomenological models must include complementary bands above and below the flat bands to accurately capture the symmetry of the electronic bands of tBLG.
@@
While our results provide a high level of detail of the structure and evolution of the flat-band wavefunctions in undoped tBLG, we note that our calculations are based on DFT, which is a single-particle approach. As such, they do not capture strong correlation effects, which play an important role in tBLG. Extending our approach to include many-body effects, finite doping, and the direct calculation of topological invariants are important directions for future work.
IV Acknowledgments
@@
The authors acknowledge support from the Simons Foundation award No. 896626 and the US Army Research Office (ARO) MURI project under Grant No. W911NF-21-0147. A.Z. is supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0025528. D.B. acknowledges support from the NTU Startup Grant (Award Number 025661-00003). M.M.E.A. acknowledges support from the Ministry of Education, Singapore (Research Centre of Excellence award to the Institute for Functional Intelligent Materials, I-FIM, project No. EDUNC-33-18-279-V12), the National Research Foundation, Singapore, under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-028). E. M. acknowledges the hospitality of the Harvard Physics Department, where part of this work was done. Computing resources were provided by the Harvard University FAS Division of Science Research Computing Group and the Texas Advanced Computing Center (TACC) at The University of Texas at Austin.